# What will be the Euclidean dimension of an Ising ferromagnetic cubic shell?

### v1

@@ 3.3 Metastable behaviour:
To study the decay of the metastable state in the Ising ferromagnetic cubic shell we have initially considered a cubic shell of size $L=40$ with all Ising spins up, kept at a fixed temperature $T=1.82=0.8T_{c}^{p}$ below pseudo-critical temperature. Now the system is kept in an external magnetic field, say $h_{ext}=-0.12$, applied in the opposite direction of that of initial magnetisation. The system is observed to keep itself in a metastable state of positive magnetisation for a certain time duration, even in the presence of a negative external magnetic field and finally decay to the stable state of negative magnetisation, as shown in figure 6(a). It may be noted here, that multiple metastable states (multiple step like variation of magnetisation) are found in the case of cubic shell. This has not been observed [9] before in the case of two dimensional Ising ferromagnet. This is an additional indication that the Ising ferromagnetic cubic shell behaves differently from that of a two dimensional Ising ferromagnet.
@@
In our simulation, $\tau_{nucleation}$ is computed as the mean value of nucleation time (or metastable lifetime) to achieve $m=0.7$ across 1000 random samples in the strong field regime (SFR) and coalescence regimes (CR) where fluctuations are comparatively small. Conversely, in the nucleation regime (NR) or single droplet regime, where fluctuations are enormous we have considered the median value of nucleation times instead of the mean. This prescription has been followed by many related previous studies[9]. Now, we have examined the nucleation time $\tau_{nucleation}$ as a function of the external field $h_{ext}$ for different sizes of cubic shell ($L=30$, 40 and 50) to analyze the impact of the external field on nucleation (or metastability). The nucleation time $\tau_{nucleation}$ is plotted against $1/|h_{ext}|$ and $1/h_{ext}^{2}$ in semi-logarithmic scale represented in figure 8 (a) and (b) respectively. In these graphical plots three distinct regimes (i) strong field regime (SFR), (ii) coalescence regime (CR) and (iii) nucleation regime (NR) are definitively characterized. It is quite noticeable that the metastable lifetime (or nucleation time) remains almost independent of system size in both strong field regime (SFR) and coalescence regime (CR). On the other hand, at weaker field namely nucleation regime(NR), the nucleation time $\tau_{nucleation}$ decreases with increasing the system size($L$) [9] for same amount of applied magnetic field $h_{ext}$.
@@
The nucleation process significantly depends on the structure, dimensionality and constraints of the lattice. In a three dimesional solid cubic lattice, each spin, inside the bulk, interacts with six nearest neighbours, while each spin at the surface (excluding the edge) can interact with five interacting spins. In contrast, spin in the cubic shell lattice can experience the interaction with a maximum of four nearest neighbours. Consequently, the metastable behaviour of the cubic shell should differ from that of a three-dimensional solid cubic. Notably, the nucleation regime of a cubic shell initiates at a weaker field compared to that of the solid cube. The results obtained for the cubic shell are shown in figure 8(b) while Fig-9 shows the results for the solid cube.
@@
Where ‘$d$’ denotes the system’s Eucleadian dimensionality, $\sigma$ stands for surface tension and $K_{d}$ is a $d$ dependant constant. In the coalescence regime, multiple supercritical droplets form and grow simultaneously, they coalesce to form bigger droplets and finally engulf the whole system. In the multi-droplet or coalescence regime, the metastable lifetime (or nucleation time) is given by, [4]
@@
In the coalescence regime (e.g $h_{ext}=-0.12$) a pair of supercritical droplets (connected domain) of down spin form and evolve simultaneously on the faces 1 and 4 of the cubic shell lattice and after a certain time iterations these two droplets merge along the edge boundary of the lattice, creating a larger domain of negative spins, as shown in figure 13. Figure 13 (a) illustrates how two clusters are created at the time after the system achieves the state of $m=0.7$. Subsequently, another two droplets of down spin, larger than critical size nucleate on face 2 and face 3 around the same time and gradually coalesce with other droplets. This process continues until the entire system (all planes) is nearly engulfed by the droplets or clusters of negative spins. The nucleation phenomena in the Ising cubic shell at Coalescence regime are clearly explained by the timelapse video of growing and shrinking droplets, found on YouTube with the link https://youtu.be/7AeoOowC07g. Whereas in the nucleation regime (e.g. $h_{ext}=-0.08$), a single droplet of supercritical size on face 5 grows to occupy the majority of that face, as demonstrated in figure 14. A similar phenomenon successively occurs on the other faces of the cubic shell lattice (e.g. 3, 4, 2 and 1 etc.). The single droplet of negative spin forms at a time on each face one after another and then gradually evolves to invade the system, as explained by the timelapse video of growing droplets, found on YouTube with the link https://youtu.be/QyNjoeC8rSg. It is worth mentioning that a single droplet (cluster) developed on one of the six faces of the cubical shell does not penetrate to other faces which confirms the growth in the nucleation regime.
@@ 4 Summary:
The equilibrium and nonequilibrium properties of an Ising ferromagnetic cubic shell have been studied extensively by Monte Carlo simulation. The results are compared with that of a solid cube. The cubical shell differs from a solid cube by the number of nearest neighbours of the spins on the surface. Geometrically, the Euclidean dimensions of both the solid cube and the cubical shell are three. However, it is noticed, from detailed Monte Carlo analysis that the cubical shell behaves like a two-dimensional object with respect to its equilibrium ferromagnetic properties of phase transition. Our results of the equilibrium magnetisation as a function of the temperature of a cubical shell show ferromagnetic phase transition near a pseudocritical temperature which is close to that for the two-dimensional Ising ferromagnet. That is an interesting result, we believe. The transition temperature of the cubical shell has been found to be the temperature which maximizes the susceptibility. The pseudocritical temperature $T^{p}_{c}\simeq 2.28$ is close to the value obtained from the exact calculation by Onsager[17].
@@
The surprising results are obtained in studying the relaxation behaviour of the Ising ferromagnetic cubic shell. Expecting the exponential relaxation (as mentioned in earlier studies of a two dimensional Ising ferromagnet), we got surprised when we have observed the non-exponential relaxation of cubic shell. Moreover, our results show that the relaxation of an Ising ferromagnetic cubic shell is not even stretched exponential[20]. In this case, the cubic shell does not behave like two dimensional system as predicted from the results of ferro-para phase transition of Ising ferromagnetic cubic shell.
@@
We have also investigated the transient behaviours of the cubical Ising ferromagnetic shell. The well-persistent multiple metastability has been observed in the weak field limit. The metastable lifetime has been calculated. This metastable lifetime has been studied as a function of the applied magnetic field. The three regions namely, the strong field regime, coalescence regime and nucleation regimes have been identified distinctly. These results are qualitatively consistent with the prediction of Becker-Döring classical nucleation theory. However, the statistical analysis matches more likely with the two-dimensional metastable behaviour of Ising ferromagnet.
@@
In conclusion, we can say that the Euclidean dimensions of an Ising ferromagnetic cubic shell cannot be determined certainly, in the case of its equilibrium and nonequilibrium properties. It would be interesting to study the variations of the ferro-para transition temperature with the thickness of the cubic shell. This requires heavy computational efforts and is beyond the scope of the present study.
@@
Authors’ contributions: Ishita Tikader developed the code, prepared the figures, and wrote the manuscript. Muktish Acharyya-Conceptualized the problem, analysed the results and wrote the manuscript.

### v2

@@ 3.3 Metastable behaviour:
To study the decay of the metastable state in the Ising ferromagnetic cubic shell we have initially considered a cubic shell of size $L=40$ with all Ising spins up, kept at a fixed temperature $T=1.82=0.8T_{c}^{p}$ below pseudo-critical temperature. Now the system is kept in an external magnetic field, say $h_{ext}=-0.12$, applied in the opposite direction of that of initial magnetisation. The system is observed to keep itself in a metastable state of positive magnetisation for a certain time duration, even in the presence of a negative external magnetic field and finally decay to the stable state of negative magnetisation, as shown in figure 6(a). It may be noted here, that multiple metastable states (multiple step like variation of magnetisation) are found in the case of cubic shell. This has not been observed [9] before in the context of two-dimensional Ising ferromagnet. This is an additional indication that the Ising ferromagnetic cubic shell behaves differently from that of a two-dimensional Ising ferromagnet.
@@
In our simulation, $\tau_{nucleation}$ is computed as the mean value of nucleation time (or metastable lifetime) to achieve $m=0.7$ across 1000 random samples in the strong field regime (SFR) and coalescence regimes (CR) where fluctuations are comparatively small. Conversely, in the nucleation regime (NR) or single droplet regime, where fluctuations are enormous we have considered the median value of nucleation times instead of the mean. This prescription has been followed by many related previous studies[9]. Now, we have examined the nucleation time $\tau_{nucleation}$ as a function of the external field $h_{ext}$ for different sizes of cubic shell ($L=30$, 40 and 50) to analyze the impact of the external field on nucleation (or metastability). The nucleation time $\tau_{nucleation}$ is plotted against $1/|h_{ext}|$ and $1/h_{ext}^{2}$ in semi-logarithmic scale represented in figure 8 (a) and (b) respectively. In these graphical plots three distinct regimes (i) strong field regime (SFR), (ii) coalescence regime (CR) and (iii) nucleation regime (NR) are definitively characterized. It is quite noticeable that the metastable lifetime (or nucleation time) remains almost independent of system size in both strong field regime (SFR) and coalescence regime (CR). On the other hand, at weaker field namely nucleation regime(NR), the nucleation time $\tau_{nucleation}$ decreases with increasing the system size($L$) (as reported in [9] for two- and three-dimensional Ising ferromagnetic system) for same amount of applied magnetic field $h_{ext}$.
@@
The nucleation process significantly depends on the structure, dimensionality and constraints of the lattice. In a three dimesional solid cubic lattice, each spin, inside the bulk, interacts with six nearest neighbours, while each spin at the surface (excluding the edge) can interact with five interacting spins. In contrast, spin in the cubic shell can experience the interaction with a maximum of four nearest neighbours. Consequently, the metastable behaviour of the cubic shell should differ from that of a three-dimensional solid cubic. Notably, the nucleation regime of a cubic shell initiates at a weaker field compared to that of the solid cube. The results obtained for the cubic shell are shown in figure 8(b) while Fig-9 shows the results for the solid cube.
@@
Where ‘$d$’ denotes the system’s Euclidean dimensionality, $\sigma$ stands for surface tension and $K_{d}$ is a $d$ dependant constant. In the coalescence regime, multiple supercritical droplets form and grow simultaneously, they coalesce to form bigger droplets and finally engulf the whole system. In the multi-droplet or coalescence regime, the metastable lifetime (or nucleation time) is given by, [4]
@@
In the coalescence regime (e.g $h_{ext}=-0.12$) a pair of supercritical droplets (connected domain) of down spin form and evolve simultaneously on the faces 1 and 4 of the cubic shell and after a certain time iterations these two droplets merge along the edge boundary of the lattice, creating a larger domain of negative spins, as shown in figure 13. Figure 13 (a) illustrates how two clusters are created at the time after the system achieves the state of $m=0.7$. Subsequently, another two droplets of down spin, larger than critical size nucleate on face 2 and face 3 around the same time and gradually coalesce with other droplets. This process continues until the entire system (all planes) is nearly engulfed by the droplets or clusters of negative spins. The nucleation phenomena in the Ising cubic shell at Coalescence regime are clearly explained by the timelapse video of growing and shrinking droplets, found on YouTube with the link https://youtu.be/7AeoOowC07g. Whereas in the nucleation regime (e.g. $h_{ext}=-0.08$), a single droplet of supercritical size on face 5 grows to occupy the majority of that face, as demonstrated in figure 14. A similar phenomenon successively occurs on the other faces of the cubic shell (e.g. 3, 4, 2 and 1 etc.). The single droplet of negative spin forms at a time on each face one after another and then gradually evolves to invade the system, as explained by the timelapse video of growing droplets, found on YouTube with the link https://youtu.be/QyNjoeC8rSg. It is worth mentioning that a single droplet (cluster) developed on one of the six faces of the cubical shell does not penetrate to other faces which confirms the growth in the nucleation regime.
@@ 4 Summary:
The equilibrium and nonequilibrium properties of an Ising ferromagnetic cubic shell have been studied extensively by Monte Carlo simulation. The results are compared with that of a solid cube. The cubical shell differs from a solid cube by the number of nearest neighbours of the spins on the surface. Geometrically, the Euclidean dimensions of both the solid cube and the cubical shell are three. However, it is noticed from detailed Monte Carlo analysis that the Ising ferromagnetic cubical shell behaves like a two-dimensional object in the context of ferro-para phase transition. Our results of the equilibrium magnetisation as a function of the temperature of a cubical shell show ferromagnetic phase transition near a pseudocritical temperature which is close to that for the two-dimensional Ising ferromagnet. That is an interesting result, we believe. The transition temperature of the cubical shell has been found to be the temperature which maximizes the susceptibility. The pseudocritical temperature $T^{p}_{c}\simeq 2.28$ is close to the value of that obtained from the exact calculation by Onsager[17].
@@
The surprising results are obtained in studying the relaxation behaviour of the Ising ferromagnetic cubic shell. Expecting the exponential relaxation (as mentioned in earlier studies of a two dimensional Ising ferromagnet), we got surprised when we have observed the non-exponential relaxation of cubic shell. Moreover, our results show that the relaxation of an Ising ferromagnetic cubic shell is not even stretched exponential[22]. In this case, the cubic shell does not behave like two-dimensional system as predicted from the results of ferro-para phase transition of Ising ferromagnetic cubic shell.
@@
We have also investigated the metastable behaviours of the cubical Ising ferromagnetic shell. The well-persistent multiple metastability has been observed in the weak field limit. The metastable lifetime has been calculated. This metastable lifetime has been studied as a function of the applied magnetic field. The three regions namely, the strong field regime, coalescence regime and nucleation regimes have been identified distinctly. These results are qualitatively consistent with the prediction of Becker-Döring classical nucleation theory. However, the statistical analysis matches more likely with the two-dimensional metastable behaviour of Ising ferromagnet.
@@
In conclusion, we can say that the Euclidean dimension of an Ising ferromagnetic cubic shell cannot be determined certainly, in the case of its equilibrium and nonequilibrium magnetic properties. It would be interesting to study the variations of the ferro-para transition temperature with the thickness of the cubic shell. This requires huge computational efforts and is beyond the scope of the present study. Moreover, the critical exponents for the magnetisation ($M(T_{c})\sim L^{-{{\beta}\over{\nu}}}$) and susceptibility ($\chi_{max}\sim L^{{\gamma}\over{\nu}}$) can be estimated (through detail finite size analysis) to capture the effects of the Euclidean dimension ($d$) of the system. These will be considered in a separate project and the results will be reported elsewhere. The metastable behaviours of other two dimensional systems (e.g., Kagome lattice) would be an interesting study.
@@
Authors’ contributions: Ishita Tikader- Developed the code, prepared the figures, and wrote the manuscript. Muktish Acharyya-Conceptualized the problem, analysed the results and wrote the manuscript.